OpenStudy (anonymous):

To begin the experiment, 1.11g of methane CH4 is burned in a bomb calorimeter containing 1000 grams of water. The initial temperature of water is 24.85C. The specific heat of water is 4.184 J/g C. The heat capacity of the calorimeter is 695 J/ C . After the reaction the final temperature of the water is 35.65C. Calculate the change in temperature, ΔT

OpenStudy (anonymous):

@Photon336 Can you help me with this

OpenStudy (photon336):

are you familiar with this formula \[q = m~C~\Delta T \] q = heat m = mass grams C = specific heat in J/gC Temperature is in Celcius wait, they are just asking for the delta T?

OpenStudy (anonymous):

ive been out sick for a couple weeks so im a bit behind but the question after that asks Using the formula water q =m•c• ΔT ,calculate the heat absorbed by the water so im guessing i use it for this as well

OpenStudy (anonymous):

im just not to sure how to apply it

OpenStudy (photon336):

Well, I'm pretty sure we would use the same formula

OpenStudy (photon336):

here is what we know \[q = m~c~\Delta~T\] 1,000 grams of water 4.184 j/gC specific heat of water and we also know delta T \[\Delta T = (T_{temperature~final}-T_{temperature~initial})\]

OpenStudy (photon336):

you would plug these values into your equation for the heat. also, Q = HEAT, either absorbed or released. if HEAT has a negative sign we know that heat is released. if it has a positive sign we know heat is absorbed.

OpenStudy (anonymous):

would the 24.85c be the q?

OpenStudy (anonymous):

oh wait i see what you were trying to say

OpenStudy (anonymous):

so i would do 35.65-24.85

OpenStudy (photon336):

@bambi152 it's better to take a look at it this way \[1,000~grams*(4.184~\frac{ joules }{ grams~Celsius })*(35.65-24.85~Celcius) = joules\]

OpenStudy (photon336):

brb

OpenStudy (anonymous):

the part thats hardest for me is changing it to the right unit

OpenStudy (photon336):

Take a look at this see what happens? this is called dimensional analysis \[1,000~ \cancel\grams*(4.184~\frac{ joules }{ \cancel\grams ~ \cancel\Celsius })*(35.65-24.85~\cancel\Celcius) = joules\]

OpenStudy (photon336):

energy is in joules right? if you notice, i'm left with joules as the only unit remaining

OpenStudy (anonymous):

so then i just calculate 1000*4.184*(35.65-24.85)

OpenStudy (anonymous):

=45187.2 ?

OpenStudy (photon336):

Joules

OpenStudy (rushwr):

Hold on @Photon336 don't we have to add the heat capacity of the calorimeter ?

OpenStudy (rushwr):

Actually the heat released by the combustion of CH4 is absorbed completely by water and the calorimeter (If we assume that no heat is released to the environment) And that assumption should be made to find the heat nah ! If we consider the heat capacity of the calorimeter as C Heat dissipated as q mass of water as 'm' specific heat capacity of water as 's' and the temperature change as 'T' q= (C+ms) T

OpenStudy (rushwr):

units are correct but the heat capacity of the calorimeter needs to be added in to the sum

OpenStudy (photon336):

Yeah, I didn't even take that into account at all

OpenStudy (photon336):

so you're adding the two heat capacities together?

OpenStudy (rushwr):

yes right ? C=ms so ms can be substituted to water's heat capacity ! Because calorimeter too absorbs heat

OpenStudy (rushwr):

\[q=(C+ms)T\] \[q= (695JC ^{-1}+ 1000g * 4.184Jg ^{-1}C ^{-1}) 10.8C\]

OpenStudy (rushwr):

So the answer will come in Joules

OpenStudy (rushwr):

@bambi152 I hope you get it :)